Sami has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. how many dimes does she have?
Answer by stanbon(75887)

Sami has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. how many dimes does she have?
--------------------------
Equations:
5n + 10d = 565 cents
5(n-8)+10(2d) = 1045 cents
----------------------------------
Rearrange:
n + 2d = 113
(n-8) + 4d = 209
-------------------------
n + 2d = 113
n + 4d = 217
--------------------
2d = 104
d = 52 (# of dimes)
---------------------------
Substitute into n + 2d = 113 and solve for "n":
n + 2(52) = 113
n = 9 (# of nickels)
==========================
Cheers,
Stan H.
==========================